### v2

# Aerial Vision-Language Navigation with a Unified Framework for Spatial, Temporal and Embodied Reasoning

Abstract
Aerial Vision-and-Language Navigation (VLN) aims to enable unmanned aerial vehicles (UAVs) to interpret natural language instructions and navigate complex urban environments using onboard visual observation. This task holds promise for real-world applications such as low-altitude inspection, search-and-rescue, and autonomous aerial delivery. Existing methods often rely on panoramic images, depth information, or odometry to support spatial reasoning and action planning. These requirements increase system cost and integration complexity, thereby hindering the practical deployment of lightweight UAVs. In this work, we present a unified aerial VLN framework that operates solely on egocentric monocular RGB observations and natural language instructions. The model formulates navigation as a next-token prediction problem, jointly optimizing spatial perception, trajectory reasoning, and action prediction through prompt-guided multi-task learning. Moreover, we propose a keyframe selection strategy to reduce visual redundancy by retaining semantically informative frames, along with an action merging and label reweighting mechanism that mitigates long-tailed supervision imbalance and facilitates stable multi-task co-training. Extensive experiments on the AerialVLN and OpenFly benchmark validate the effectiveness of our method. Under the challenging monocular RGB-only setting, our model achieves strong results across both seen and unseen environments. It significantly outperforms existing monocular baselines and narrows the performance gap with state-of-the-art panoramic RGB-D counterparts. Comprehensive ablation studies further validate the contribution of our architectural design and choices.
I Introduction
Unmanned Aerial Vehicle (UAV) has become an indispensable tool in modern remote sensing applications, playing a central role in infrastructure inspection, environmental monitoring, and emergency response [17]. Previous research has largely focused on passive perception tasks, including object detection [21, 9] and tracking [7, 35] from aerial images or videos, without interaction with the world. In contrast, aerial navigation tasks require the drone to perceive, reason, and act in dynamic environments. Recently, aerial Vision-and-Language Navigation (VLN) [25] has emerged as a new paradigm, where drones follow high-level language instructions to navigate the destination through 3D outdoor environments. By leveraging natural language as a human-centric interface, aerial VLN significantly reduces the reliance on expert pilots, lowers the barrier of human-UAV interaction, and enables intuitive task specification in high-stakes scenarios.
Recent works have explored the design of benchmarks and datasets to facilitate research in aerial vision-and-language navigation. AVDN [11] first proposed a dialogue-based setting involving asynchronous interactions between a human commander and a UAV agent. AerialVLN [25] introduced high-fidelity city-scale simulations with diverse human-annotated trajectories. CityNav[20] extended this line by leveraging real-world urban reconstructions. OpenUAV [33] formulated aerial VLN as full-trajectory prediction with human-in-the-loop evaluation, and OpenFly [12] scaled up scene and instruction diversity via automatic generation. AeroDuo [36] explores the collaborative setting with multi-agent instruction following with altitude-aware role assignment. Moreover, CityEQA [49] and 3D Open-EQA [46] have proposed embodied question answering benchmarks to assess model’s perception and reasoning capabilities from an aerial perspective. However, as illustrated in Fig. 1, aerial VLN requires an agent to continuously integrate current observation, trajectory history, and high-level instruction. Concretely, the agent must recognize location from egocentric views, assess task progress relative to instruction, and infer the next action consistent with the intended route. Aerial VLN presents unique challenges across three core aspects.
Figure 1: Aerial vision-language navigation. Left: A drone receives a natural-language instruction along with egocentric visual observations and is required to navigate to the destination in a complex outdoor environment. Right: This task relies on the agent’s ability to maintain an accurate understanding of its navigational situation, including estimating its current position, interpreting its progress within the instruction, and determining the next movement consistent with the described route. The example highlights these dimensions of temporal and spatial reasoning, which are central to reliable long-horizon aerial navigation.
  1. 1.
    Unified Vision-Language-Action Alignment: UAVs struggle for high-dimensional action spaces involving both horizontal and vertical movements, where altitude changes affect visibility, orientation, and control feasibility. Mapping free-form natural language instructions to executable flight commands requires reliable alignment between egocentric semantics and spatial geometry.
  2. 2.
    Large-scale Outdoor Aerial Scenes: UAVs operate in large-scale urban environments with dense infrastructures, complex spatial layouts, and highly dynamic semantic distributions. This complexity demands precise spatial perception (”ascend to the height of the street lamp”) and the ability to robustly ground language-referred landmarks (”the gray house with a slope”) using egocentric observations.
  3. 3.
    Long-Horizon Temporal Reasoning: Aerial instructions often describe multi-stage flight plans that span long distances and evolving visual contexts. Agents must not only interpret sequential goals but also maintain trajectory-level awareness, track historical behavior, and align current decisions with global navigation intent.
Massive endeavors have attempted to alleviate these challenges. One line of research leverages pre-trained vision language models (VLMs) to enable zero-shot instruction following [13, 45, 39]. These approaches introduce explicit intermediate representations and memory mechanisms to mitigate large-scale spatial reasoning difficulties and preserve long-term spatial context. A large language models are subsequently employed for high-level planning. However, such modular designs decouple perception, representation, and decision-making into separate stages, which may introduce error accumulation and restrict joint optimization across complementary navigation capabilities. Another line of research adopts learning-based end-to-end modeling to directly map visual observations and language instructions to embodied action sequences [25, 48, 20]. These methods enhance cross-modal alignment by incorporating geographic spatial representations or history-aware fusion modules to better utilize spatial cues and history observations. More recent efforts further leverage large vision-language models with strong multimodal capabilities to improve reasoning over aerial trajectories [12, 5, 38]. While these approaches have achieved promising performance, many rely on auxiliary inputs such as panoramic cameras, depth sensors, or odometry, which increase hardware complexity and deployment cost. Moreover, the limited memory context of current models remains a bottleneck, making long-horizon aerial navigation particularly challenging.
To address the aforementioned limitations without introducing additional sensing modalities, while leveraging the strong multimodal capability of vision-language models, we develop a unified framework for aerial VLN that relies solely on onboard egocentric RGB observations and natural language instructions. The proposed framework directly predicts the next navigation action, bridging language semantics with visual context. we formulate navigation under a Next-Token Prediction (NTP) paradigm, jointly modeling spatial perception, trajectory reasoning, and action generation within a single vision-language backbone. This unified formulation enables cross-modal alignment and trajectory-level reasoning to be optimized under a shared objective without architectural modification. Within this framework, Task-specific prompts are introduced to steer learning towards distinct reasoning objectives while preserving architectural simplicity. Furthermore, considering the structural characteristics of aerial navigation, including substantial visual redundancy during continuous flight and imbalance in action distributions, we incorporate an aerial-specific design consisting of keyframe selection, action merging, and label reweighting. These components are integrated into the overall learning framework to improve supervision balance and sequence modeling stability. Experimental results across multiple benchmarks demonstrate that our method consistently outperforms baseline models.
In summary, our paper makes the following contributions:
  1. 1.
    Unified Next-Token Formulation: We formulate aerial VLN as a next-token prediction problem, jointly modeling spatial perception, trajectory reasoning, and action generation within a single autoregressive backbone, thereby facilitating tighter cross-modal alignment.
  2. 2.
    Prompt-Driven Multi-Task Supervision: We introduce task-specific prompts to incorporate auxiliary objectives for spatial grounding and trajectory reasoning, strengthening the model’s reasoning ability over spatial structure and temporal evolution.
  3. 3.
    Aerial-Specific Design: We propose an aerial-specific design and training strategy, including keyframe selection, action merging, and label reweighting, to address visual redundancy and action imbalance inherent in aerial navigation for stable training.
II Related Works
II-A Vision-and-Language Navigation
Vision-and-language navigation (VLN) tasks require an agent to follow natural language instructions using visual observations to achieve a target location [37, 14]. Early VLN benchmarks formulated navigation on a predefined graph of panoramic viewpoints, where the agent selects discrete actions to move between connected nodes [3, 19, 42]. To move beyond the discrete setting, VLN-CE [18] replaced the navigation graph with free movement in simulation, where the agent executes mid-level actions, such as moving forward and turning. A variety of learning-based methods have been proposed under this setting, including cross-modal attention [15], memory-augmented agents [34], and planning-based approaches [2, 30, 6]. A growing line of work has focused on training vision-language models that directly map RGB observations and instructions to navigation actions. NaVid [44] demonstrates that a vision-language model trained on large-scale continuous trajectories can effectively use monocular RGB video to model temporal context and predict actions, achieving strong performance. NaVILA [10] introduces a two-stage framework that maps visual observations and instructions into mid-level language actions, which are then executed by a learned locomotion policy, enabling instruction-conditioned control in low-level continuous environments. MonoDream [32] introduces auxiliary representation learning tasks that train a monocular agent to align visual semantics and action intent by predicting latent panoramic features. Uni-NaVid [43] unifies diverse navigation tasks within a single end-to-end framework by learning from multi-task RGB video data. VLN-R1 [28] combines supervised and reinforcement fine-tuning to enhance temporal reasoning and multi-step action planning.
While these approaches have made substantial progress, they are primarily developed for ground-based agents in indoor environments and typically operate within limited spatial scales. Our work instead focuses on aerial VLN in large-scale urban environment, with inherent challenges, such as viewpoint variability, long-range flight paths, and complex 3D outdoor scenes. we adopt a unified framework and enhance it with prompt-guided auxiliary supervision to strengthen spatial perception and trajectory reasoning over long horizons
II-B Aerial Vision-and-Language Navigation
Aerial navigation has attracted growing attention, with early efforts focusing on conventional tasks such as obstacle avoidance [23, 47] and object tracking [31, 29] based on onboard cameras or GPS signals. Recent advances explore language-guided aerial navigation, aiming to empower UAVs to interpret and execute natural language instructions in large-scale, open-world scenarios [25]. This task introduces additional challenges beyond visual navigation, such as grounding language instructions into visual perception and continuous control, performing temporal reasoning, and aligning multimodal information for robust decision-making.
Research on aerial VLN has gradually evolved with the introduction of dedicated benchmarks and diverse solution paradigms. AVDN [11] introduces a dialogue-based UAV-VLN setting while AerialVLN [25] and CityNav [20] construct large-scale outdoor datasets using synthetic cities and real-world 3D reconstruction, respectively. OpenUAV [33] formulates UAV-VLN as full-trajectory prediction and incorporates human-in-the-loop evaluation. OpenFly [12] scales up dataset diversity through automatic generation over varied high-fidelity virtual environments. Early approaches to aerial VLN adapted models from ground-based navigation, but often struggled with dynamic viewpoints and large-scale outdoor complexity. More recently, with the rise of large language models (LLMs) and vision language models (VLMs), a line of work explores training-free paradigms for zero-shot aerial VLN. These methods convert visual observations into text prompts in the form of topological graphs, symbolic maps, or structured semantic matrices and rely on LLMs to infer the next navigation action or waypoint. Among them, STMR [13] introduces a Semantic-Topo-Metric Representation that encodes current view, pose, and subgoal cues into matrix-form prompts for spatial reasoning over landmarks and directions. CityNavAgent [45] further enhances semantic grounding via symbolic 2D landmark maps derived from real aerial imagery and spatial detection. In contrast, learning-based approaches directly align visual observations and language instructions through end-to-end optimization. LAG [25] introduces a look-ahead guidance mechanism to Cross-Modal Attention method [18]. Grid-based View Selection [48] reformulates aerial VLN as a grid-based view selection problem with coupled vertical-horizontal action modeling, bird’s-eye history fusion, and cross-modal alignment. More closely related to our work, OpenFly [12] proposes keyframe-aware modeling and adaptive token merging to improve efficiency and stability in long-horizon navigation.
However, many existing aerial VLN approaches rely on additional sensory inputs or explicit structural representations to facilitate navigation. Differently, we focus on an egocentric monocular RGB setting and formulate aerial VLN under a unified next-token prediction paradigm. Within this formulation, we incorporate two auxiliary objectives to enhance the scene understanding and temporal reasoning. We further introduce aerial-specific training strategies, including keyframe selection, action merging, and label reweighting, to reduce visual redundancy and mitigate label imbalance.
Figure 2: Trajectory statistics before and after data preprocessing, showing that action merging and keyframe selection yield a richer action space and more compact navigation sequences. For brevity, we use MF for move forward, etc.
III Method
III-A Task Definition
Aerial vision-and-language navigation. The goal of aerial VLN is to learn a policy that enables drones to follow high-level language instructions to navigate through 3D outdoor environments and reach the target location. Specially, an aerial agent is tasked with navigating a continuous 3D space according to a natural language instruction $I=\{w_{1},w_{2},\ldots,w_{N_{I}}\}$, which comprises $N_{I}$ words describing the desired flight route. At each time step $t$, the agent receives a visual trajectory $\tau_{t}=\{x_{1},\ldots,x_{t}\}$, where $x_{t}\in\mathbb{R}^{H\times W\times 3}$ is the first-person RGB image captured by its onboard camera. Conditioned on visual trajectory $\tau_{t}$ and the instruction $I$, the agent selects an action $a_{t}$ from a discrete action set $\mathcal{A}$. Executing $a_{t}$ moves the agent to the next state in the environment, yielding the subsequent image observation $x_{t+1}$. An episode terminates when the agent chooses the stop action or reaches the maximum step limit. A navigation episode is considered successful if the agent’s final position lies within $20$ meters of thedestination. The action space and corresponding motion magnitudes are defined according to the specific environment settings. In AerialVLN environment, $\mathcal{A}=\{\textit{move forward},\allowbreak\textit{turn left},\allowbreak\textit{turn right},\allowbreak\textit{ascend},\allowbreak\textit{descend},\allowbreak\textit{move left},\allowbreak\textit{move right},\allowbreak\textit{stop}\}$. horizontal motions displace the agent by $5$ units, vertical motions by $2$ units, and turning applies a fixed $15^{\circ}$ change in orientation. In OpenFly environment, $\mathcal{A}=\{\textit{move forward},\allowbreak\textit{turn left},\allowbreak\textit{turn right},\allowbreak\textit{ascend},\allowbreak\textit{descend},\allowbreak\textit{ stop}\}$. Horizontal motions displace the agent by $3$ units, vertical motions by $3$ units, and turning applies $30^{\circ}$.
Figure 3: Overview of our framework. Given egocentric keyframes selected from the onboard video stream, our model first encodes the visual observations through a vision encoder and an MLP projector to obtain visual tokens, while language instructions are processed by a text tokenizer. The unified multimodal tokens are then fed into a large language model that is jointly trained on three complementary tasks: (i) Spatial Perception, which queries the current scene; (ii) Trajectory Reasoning, which summarizes historical motion and infers the agent’s navigational context; and (iii) Embodied Navigation, which predicts high-level action commands. The predicted textual action is further parsed and decomposed into a sequence of predefined motion primitives for execution in the physical environment.
III-B Data Preprocessing
In data preprocessing stage, we refine the raw drone trajectories by applying action merging and keyframe selection, two complementary steps that improve the temporal structure of training samples. Together, these preprocessing steps organize trajectories into clearer motion segments paired with representative visual cues, improving the quality of supervision for navigation learning. As shown in Fig. 2, our approach expands the action vocabulary, produces a more balanced action distribution, and significantly shortens trajectories by eliminating redundant micro-actions.
Action Merging. UAV navigation trajectories are typically long and dominated by frequent forward micro-steps, leading to highly imbalanced and fragmented action sequences. To reduce fragmented micro-actions and enhance the semantic clarity of motion patterns, we merge consecutive identical actions into a single segment with a bounded length (e.g., combining three turn left $15^{\circ}$ steps into a single turn left $45^{\circ}$ step). This produces more meaningful action units and yields a more balanced and diverse action distribution.
Keyframe Selection. To further suppress temporal visual redundancy, we extract keyframes at the boundaries of the merged action segments. This simple yet effective strategy is motivated by the observation that abrupt changes in the agent’s motion—such as transitioning from straight movement to a turning action—are often caused by the appearance of landmarks. These turning points naturally delimit semantically meaningful transitions in the underlying trajectory. By selecting the frames at these boundaries as keyframes, we obtain a compact set of observations that preserves landmarks and other salient visual cues while filtering out redundant intermediate views. This yields a more informative and temporally structured visual stream for downstream navigation learning. Unlike simple uniform sampling [10] or detector-dependent landmark selection [12], our strategy is control-aligned and does not rely on auxiliary perception modules, providing an efficient keyframe selection mechanism.
III-C Framework Overview
As illustrated in Fig. 3, our framework tackles the aerial vision-and-language navigation task by processing natural-language instructions together with an egocentric video stream under a unified next-token prediction paradigm.
Overall Pipeline. At each timestep, we sample a compact set of keyframes from the observation trajectory so far and pair them with the instruction. The keyframes are encoded into visual tokens through a vision encoder and projected into the language embedding space, while the instruction is tokenized independently. The resulting multimodal token sequence is fed into a large language model (LLM), which produces the next action directly in text form through autoregressive decoding. This action text is parsed into a predefined action set of low-level control commands and executed to update he drone’s state in 3D aerial environments. To strengthen the agent’s spatial understanding and its ability to interpret navigational progress, we introduce two auxiliary tasks, termed spatial perception and trajectory reasoning. The spatial perception task guides the agent to answer egocentric questions about the current scene, whereas the trajectory reasoning task requires summarizing the observation trajectory. Together, these tasks provide complementary supervision that refines the model’s representation of spatial structure and navigation dynamics, ultimately boosting aerial navigation performance.
Figure 4: Schematic diagram of the STC module with grid size $g$ = 2.
Observation Encoding and Multimodal Tokenization. At each timestep $t$, the agent receives the current egocentric image $x_{t}$ together with the accumulated visual trajectory. To obtain a compact yet informative representation, we uniformly sample $K$ frames while always retaining the first frame and the current frame. The first frame provides a stable reference of the starting viewpoint, whereas the current frame serves as the primary basis for predicting the next navigation action; the remaining historical frames supply contextual cues for interpreting navigation progress. Let the sampled keyframes be $\mathcal{X}_{t}=\{x_{1},x_{t_{2}},\ldots,x_{t_{K}}\}$ with $x_{t_{K}}=x_{t}$. Each frame is encoded by a vision backbone into patch-level features:
$E^{\text{vis}}_{t_{i}}=\mathrm{Enc}_{\text{vis}}(x_{t_{i}})\in\mathbb{R}^{N\times C}.$ (1)
Long-horizon trajectories may yield prohibitively long visual token sequences. To reduce the length of token sequences while preserving local spatial context, we introduce a Spatial Token Compression (STC) module, as illustrated in Fig. 4. Given patch-level visual tokens $E^{\text{vis}}_{t_{i}}$, STC first reshapes the token sequence into a 2D feature map. The feature map is then partitioned into non-overlapping $g\times g$ grids, where $g$ denotes the grid size. When the spatial resolution is not divisible by $g$, zero-padding is applied along the corresponding dimensions before grid partitioning. Within each grid, the $g\times g$ tokens are concatenated along the channel dimension and flattened to form a compact visual representation. This yields a compressed token sequence:
$\color[rgb]{0,0,0}\tilde{E}^{\text{vis}}_{t_{i}}=\mathrm{STC}(E^{\text{vis}}_{t_{i}})\in\mathbb{R}^{(N/g^{2})\times(Cg^{2})}.$ (2)
A lightweight MLP projector aligns visual features to the LLM embedding space:
$Z^{\text{vis}}_{t_{i}}=\mathrm{Proj}_{\text{mlp}}(\tilde{E}^{\text{vis}}_{t_{i}})\in\mathbb{R}^{(N/g^{2})\times D}.$ (3)
The navigation instruction $I=\{\omega_{1},\ldots,\omega_{L}\}$ is tokenized by the LLM tokenizer to obtain text token $Z^{\text{text}}\in\mathbb{R}^{L\times D}$. We then concatenate visual and textual tokens into a unified multimodal sequence:
$\color[rgb]{0,0,0}S_{t}=[Z^{\text{vis}}_{1},\,\ldots,\,Z^{\text{vis}}_{t_{K}},\,Z^{\text{text}}],$ (4)
which serves as the observation representation for navigation. The sequence is fed into the LLM backbone and finally we get the hidden state:
$H_{t}=\mathrm{LLM}(S_{t}),$ (5)
from which the next action $a_{t}$ is autoregressively generated in linguistic form.
Figure 5: Unified prompting interface for the proposed model. Through task-specific prompts, the model supports aerial navigation as the primary task, while also handling spatial perception and trajectory reasoning as auxiliary capabilities that enrich spatial understanding and temporal grounding.
Unified Prompting Interface. Aerial VLN poses substantially higher demands on scene understanding and temporal reasoning due to the semantic complexity of outdoor 3D environments and the long-horizon navigation trajectories. Effective navigation requires the agent to interpret fine-grained spatial cues, maintain awareness of its progression along the trajectory, and ground language instructions to visual observations across a large spatial scale. To support the acquisition of these complementary reasoning abilities, we adopt a unified prompting interface that formulates all tasks within the same next-token prediction paradigm, without any modification to the underlying model architecture. As shown in Fig. 5, we instantiate three prompt-driven task formulations that expose different levels of navigation-relevant reasoning: Spatial Perception, Trajectory Reasoning, and Embodied Navigation. Spatial perception prompts require the model to answer egocentric, scene-centric questions based on the current view (e.g., “What object appears on the right side of the image?”), thereby reinforcing fine-grained geometric and semantic grounding. Trajectory reasoning prompts ask the model to summarize the observed image sequence, encouraging structured understanding of navigation progress. Embodied navigation prompts instruct the model to predict the next control step in linguistic form, as we mentioned before. These task formulations provide complementary supervisory signals that strengthen the model’s multi-level reasoning over spatial structure and temporal trajectory evolution, thereby supporting more reliable navigation decision-making.
Action Parsing and Execution. During inference, the model operates with the embodied navigation-specific prompt as illustrated in Fig. 5, and directly generates textual navigation actions. Once the model outputs a textual action command (e.g., “The next action is move forward 15 units”), we apply a regular-expression parser to extract the action type and its numeric argument. The parsed action is then decomposed into low-level motion primitives (e.g., multiple fixed-step forward movements). These primitives are executed in an open-loop manner until all of them are completed, after which the agent receives the next visual observation from the environment, updates its trajectory history, and predicts the subsequent action. We quantitatively evaluated the reliability of this parsing module across simulated benchmarks, where it achieved a 100% success rate in extracting valid actions. This design follows a robust and established paradigm in vision-language navigation research, with prior works such as NaVid [44] and NaVid-4D [24] reporting equally reliable performance using the same approach in both simulation and real-world deployment.
III-D Training Objective
Our model is trained in a multi-task setting that jointly supervises aerial navigation, spatial perception, and trajectory reasoning through autoregressive language modeling. Since navigation trajectories exhibit a highly imbalanced distribution of actions, we apply a frequency-based label reweighting strategy. Each action $a\in\mathcal{A}$ is assigned a normalized inverse-frequency weight.
$w(a)=\sqrt{\frac{1/p(a)}{\frac{1}{|\mathcal{A}|}\sum_{b\in\mathcal{A}}1/p(b)}}.$ (6)
The overall multi-task objective integrates supervision from all three tasks. While $\mathcal{B}$ denotes the mixed training set, the loss is defined as
$\mathcal{L}$ $=\frac{1}{|\mathcal{B}|}\sum_{u\in\mathcal{B}}W(u)\,\ell_{\mathrm{CE}}(y_{u},z_{u}),$ (7)
[…]
$\mathcal{B}$ $=\mathcal{D}_{\text{nav}}\cup\mathcal{D}_{\text{sp}}\cup\mathcal{D}_{\text{tr}}.$
where $W(u)$ denotes the task-specific weight assigned to sample $u$. The term $\ell_{\mathrm{CE}}(y_{u},z_{u})$ is the standard autoregressive cross-entropy loss computed between ground truth $y_{u}$ and model output $z_{u}$. For samples from other auxiliary tasks, we apply task-specific weighting to regulate their relative influence during optimization. $\lambda_{\text{sp}}$ and $\lambda_{\text{tr}}$ are the hyperparameters of weight for spatial perception and trajectory reasoning samples, respectively.
TABLE I: Overview of the curated datasets for multi-task training.
Task Type Data Source Description
Spatial Perception Open3D-VQA (Sim env) + GQA (Real env) Drone egocentric spatial reasoning + spatial-relation QA pairs
Trajectory Reasoning AerialVLN (Sim env) Sub-trajectory segmentation + progress summaries
Aerial Navigation AerialVLN (Sim env) Step-wise tuples $(\tau_{t},I,a_{t})$ after action merging and keyframe selection
III-E Implementation Details
Dataset Curation. To support the multi-task training objectives, we curate a mixed dataset that provides complementary supervision for action prediction, spatial grounding, and temporal reasoning, as shown in Table I. Our curation is guided by two principles: maintaining a coherent input-output structure across tasks and constructing a balanced mixture that captures the diversity of aerial navigation scenarios. For aerial navigation task, we derive stepwise supervision from AerialVLN simulation trajectories by applying action merging and keyframe selection, producing compact training tuples $(\tau_{t},I,a_{t})$ . For spatial perception, we incorporate VQA-style supervision from the drone-centric subset of Open3D-VQA  [46], complemented by a spatial-relation-focused subset of GQA [16] filtered for geometric predicates. To provide temporal abstraction signals, trajectory reasoning supervision is constructed by manually segmenting training trajectories into ordered sub-trajectories (41̃6 keyframes after action merging). The segmentation follows two principles: each sub-trajectory must correspond to a semantically coherent sub-instruction, and the sub-instruction must clearly describe its terminal state or landmark. Sub-instructions are directly extracted from the original instruction without paraphrasing or language model generation to preserves the original linguistic distribution. After filtering severely ambiguous instructions, approximately $3k$ trajectories are annotated, yielding around $10k$ trajectory reasoning training instances through compositional augmentation. This curated mixture forms a coherent and comprehensive data blend for supervised fine-tuning of the proposed aerial VLN model.
Model Training. We fine-tune NVILA-lite-2B and NVILA-lite-8B [26] on our aerial navigation task, initializing both models from their pretrained weights. Following NVILA, the vision encoder is implemented using SigLIP [41], the projector is a lightweight two-layer MLP that aligns visual features with the language embedding space, and the language model is based on Qwen2 [40]. Specifically, we adopt a full fine-tuning strategy, where all model components, including the vision encoder, MLP projector, and large language model, are jointly optimized, with all parameters being trainable during training. We perform supervised fine-tuning (SFT) for our tasks and the training objective is defined by the loss function presented in Eq. 7. Training is performed on 4 NVIDIA A100 GPUs using DeepSpeed ZeRO-3 optimization, with a learning rate of $2\times 10^{-5}$. We employ a cosine decay schedule with a warm-up ratio of 0.03. We adopt a global batch size of 128, achieved through a per-GPU batch size of 8 and 4 gradient accumulation steps. All models are trained for 1 epoch. For action merging, consecutive identical actions are combined, with the maximum merge length limited to 3 steps. For keyframe sampling, we uniformly sample $K=8$ frames from the visual trajectory, while ensuring that the first and the current frames are always preserved. We set the auxiliary task weights to $\lambda_{\text{sp}}=1.0$ for spatial perception and $\lambda_{\text{tr}}=0.5$ for trajectory reasoning.
IV Experiments
IV-A Experimental Setup
Simulated environments. We conduct experiments across multiple aerial VLN benchmarks for comprehensive evaluation. AerialVLN-S benchmark [25] provides a realistic aerial navigation setting rendered in AirSim and Unreal Engine 4. The dataset covers 17 compact outdoor scenes representing diverse urban environments such as residential areas, industrial zones, parks, and villages. For evaluation, the dataset allocates 12 environments to the training and validation-seen splits, while the validation-unseen split is drawn from 5 distinct scenes. Flight trajectories are collected by expert pilots, who follow naturalistic aerial routes to ensure realistic motion patterns and diverse path configurations. Overall, AerialVLN-S provides 10,113 language instructions for training, with 333 and 531 instructions assigned to the validation-seen and validation-unseen splits, respectively. In addition to AerialVLN-S, we evaluate our method on the OpenFly benchmark [12], constructed via automatic trajectory generation and vision-language-model-based instruction synthesis. OpenFly features diverse urban environments with varied scene layouts, flight heights, and trajectory lengths. Given the large scale of OpenFly, which contains over 100k navigation trajectories, we conduct experiments on a selected subset for practical evaluation. Specifically, we use trajectories collected from six high-quality UE4 scenes, forming a subset referred to as OpenFly-S. Training and evaluation are conducted on subsets comprising 61,004 and 1,210 trajectories, respectively.
Evaluation Metrics. We adopt several standard metrics to evaluate aerial vision-and-language navigation[25]:
  • Navigation Error (NE/$m$). Measuring goal-reaching accuracy by computing the Euclidean distance between the UAV’s final position and the destination.
[…]
Baselines. We compare our approach against previous diverse baselines. These baselines fall into three categories [45]:
  • Statistical-based methods. Random Sampling uniformly samples actions from the action space while Action Sampling selects actions according to the training-set action distribution.
  • Zero-shot LLM-based methods. MapGPT [8] introduces a language-based online map representation to equip GPT-4v [1] with global spatial understanding, enabling adaptive multi-step planning for indoor VLN in a zero-shot setting. STMR[13] enhances spatial reasoning by introducing a Semantic-Topo-Metric Representation that allows LLMs to infer navigation decisions directly from structured scene information. CityNavAgent[45] performs hierarchical semantic planning to predict waypoints in a zero-shot manner, supported by a global memory module that records past trajectories for improved long-horizon navigation.
  • Learning-based methods. LingUNet [27] is an instruction-conditioned visual model adapted from goal-visible aerial navigation to step-wise VLN in AerialVLN[25] Seq2Seq [4] predicts actions using a recurrent policy that encodes visual observations and language instructions into a unified hidden representation. CMA [18] introduces cross-modal attention to better fuse RGB, depth, and textual cues for action prediction. LAG[25] adds a look-ahead guidance module. Grid-VS[48] formulates aerial VLN as a grid-based view selection task and leverages a bird’s-eye-view representation with a cross-modal transformer to model long-term navigation context and instruction alignment. NaVid [44] is a video-based vision-language model that has demonstrated strong performance in indoor VLN. OpenFly [12] adopts a keyframe-aware VLN architecture and uses adaptive token sampling over historical observations.
TABLE II: Comparison of different methods on the AerialVLN-S dataset. Observations include single RGB (S.RGB), depth (Depth), panoramic view (Pano.), and odometry information (Odo.)
Method Observation Validation Seen Validation Unseen
[…]
Grid-based VS [48] $\checkmark$ $\checkmark$ 70.3 20.8 33.4 10.2 121.3 7.4 16.1 2.5
CityNavAgent [45] $\checkmark$ $\checkmark$ $\checkmark$ 80.8 13.9 30.2 5.1 60.2 11.7 35.2 5.0
Random Sampling 109.6 0.0 0.0 0.0 149.7 0.0 0.0 0.0
Action Sampling 213.8 0.9 5.7 0.3 237.6 0.2 1.1 0.1
LingUNet [27] $\checkmark$ $\checkmark$ 383.8 0.6 6.9 0.2 368.4 0.4 3.6 0.9
Seq2Seq [4] $\checkmark$ $\checkmark$ 146.0 4.8 19.8 1.6 218.9 2.3 11.7 0.7
CMA [18] $\checkmark$ $\checkmark$ 121.0 3.0 23.2 0.6 172.1 3.2 16.0 1.1
LAG [25] $\checkmark$ $\checkmark$ 90.2 7.2 15.7 2.4 127.9 5.1 10.5 1.4
STMR [13] $\checkmark$ $\checkmark$ 96.3 12.6 31.6 2.7 119.5 10.8 23.0 1.9
MapGPT [8] $\checkmark$ 124.9 2.1 4.7 0.8 107.0 0.0 0.0 0.0
Navid [12] $\checkmark$ 105.1 6.8 15.5 1.1 106.9 6.1 12.3 0.7
OpenFly [44] $\checkmark$ 127.2 8.1 21.8 1.6 113.8 7.6 18.2 1.5
Ours $\checkmark$ 79.6 11.4 37.7 6.3 95.8 8.1 28.9 2.2
Figure 6: Qualitative comparison between predicted and ground-truth drone trajectories in 3D space. Across multiple validation episodes, the predicted paths follow the global route structure, capturing major turns, long-range transitions, and altitude changes required by language instruction. These results highlight strong 3D spatial reasoning and stable flight control from egocentric observations.
IV-B Results on Benchmarks
IV-B1 Quantitative Results on AerialVLN-S Benchmark
To ensure a fair comparison, we categorize methods by their sensor configurations, distinguishing RGB-only agents from those that rely on panoramic inputs, depth sensing, or odometry in Table II. Within the RGB-only group, our method achieves strong performance on all metrics and attains the best overall results, comparable to STMR [13], which required external depth information. The proposed method is also competitive with panoramic baselines, especially in SDTW. This is particularly meaningful in the AerialVLN setting, where many trajectories contain turn-backs and repeated segments, making SDTW a more informative measure of instruction following than SR alone. Panoramic methods typically require multiple cameras or repeated in-place rotations to build 360° observations, which increases time and energy costs for UAVs, whereas our model operates from a single-view RGB stream and thus better matches realistic deployment constraints. MapGPT [8] and Navid [44], which are strong indoor VLN baselines, perform notably worse in outdoor aerial settings, underscoring the additional challenges posed by large-scale, sparsely structured environments. Compared with the most related OpenFly [12] framework, our approach further benefits from long-horizon history modeling and auxiliary tasks, leading to more stable long-range behavior. Overall, these results establish a new state of the art for monocular RGB-only aerial VLN and validate the effectiveness of our unified framework for spatial, temporal, and embodied reasoning. Fig. 6 compares the predicted 3D flight paths with the reference routes. The predicted trajectories preserve the overall geometric structure of the routes, including the major turning segments, long-range direction changes, and the characteristic height variations implied by the instruction. Although small local deviations occasionally appear, the agent often recovers from these offsets and continues to produce geometrically coherent paths over extended horizons. The trajectories also converge toward the intended final region, reflecting stable motion patterns derived solely from egocentric RGB observations in large-scale outdoor environments.
TABLE III: Comparison of different methods on the OpenFly-S test-seen split.
Method Observation OpenFly-S Test Seen
[…]
Random Sampling 242 0.7 0.8 0.0
Seq2Seq [4] $\checkmark$ $\checkmark$ 291 0.1 3.2 0.6
CMA [18] $\checkmark$ $\checkmark$ 144 5.7 30.0 5.1
LAG [25] $\checkmark$ $\checkmark$ 130 6.6 40.8 7.0
Navid [44] $\checkmark$ 142 9.9 24.3 3.9
OpenFly [12] $\checkmark$ 98 33.2 63.5 22.8
Ours $\checkmark$ 38 54.5 72.1 49.9
Figure 7: Comparison between OpenFly and our method across three difficulty levels (Easy, Moderate, and Hard), grouped by trajectory length.
IV-B2 Quantitative Results on OpenFly-S Benchmark
We report quantitative results on the OpenFly-S benchmark in Table III. Our method achieves superior performance compared to baselines across all evaluation metrics, including NE, SR, OSR, and SPL. We further analyze success rates across different difficulty levels defined by trajectory length, following the evaluation protocol adopted in OpenFly [12]. Specifically, episodes with fewer than 30 actions are categorized as Easy, those with 30-60 actions as Moderate, and those with more than 60 actions as Hard. As shown in Fig. 7, the proposed method beats OpenFly[12] across all difficulty levels, with a larger margin on hard episodes. Notably, performance drops for both methods as the trajectory length increases, suggesting that long-horizon aerial navigation remains challenging. We observe that overall success rates on OpenFly-S are higher than those on AerialVLN-S. This behavior is consistent with the characteristics of the OpenFly dataset, which features shorter navigation trajectories and fewer action steps on average. Overall, the experimental results on both AerialVLN-S and OpenFly-S demonstrate that the proposed method maintains strong and consistent performance across multiple aerial VLN datasets constructed under different design principles, supporting its robustness to variations in trajectory scale, instruction length, and data generation strategies.
Figure 8: Visualizations of our method on AerialVLN-S benchmark. Our model successfully follows detailed long-horizon instructions, grounds visual landmarks, and executes correct spatial maneuvers across complex outdoor scenes, demonstrating strong instruction-following consistency and robust navigation ability. For clarity, colored elements denote landmarks and action phrases. Trajectories are visually grouped to facilitate analysis of long-horizon behaviors.
IV-B3 Cross-dataset Evaluation
Cross-dataset generalization is a critical capability for vision-and-language navigation models. Inspired by the evaluation protocol in ground VLN [44, 10, 32], we assess cross-dataset generalization in aerial VLN by training models on AerialVLN-S and directly evaluating them on AerialVLN without any fine-tuning. Compared to AerialVLN-S, AerialVLN presents a more challenging navigation scenario, with longer path lengths, larger-scale environments, and more complex long-horizon action sequences [25]. As Table IV shown, while all methods exhibit performance degradation under this distribution shift, our method achieves consistently better results across all metrics. These results demonstrate that our method exhibits stronger transferability and generalization capability, even under challenging cross-dataset conditions.We further report results of the proposed model trained without auxiliary tasks. Removing the auxiliary tasks results in inferior performance in the cross-dataset setting, which suggests that these tasks help enhance the agent’s spatial and temporal reasoning over the environment, thereby improving robustness under distributional shifts.
TABLE IV: Cross-dataset Generalization Evaluation on AerialVLN Val-Unseen split. Note that $\dagger$ indicates our method without auxiliary tasks.
Method Observation AerialVLN Validation Unseen
[…]
Seq2Seq [4] $\checkmark$ $\checkmark$ 352.7 0.6 4.2 0.3
CMA [18] $\checkmark$ $\checkmark$ 337.0 1.0 6.7 0.4
Ours $\checkmark$ 311.9 1.8 12.1 0.5
Ours $\checkmark$ 331.5 2.8 13.3 0.8
Figure 9: Qualitative results on auxiliary VQA tasks. Top: examples from the Open3D-VQA dataset [46], where the model answers 3D spatial queries including relative relations and geometric orientation. Bottom: examples from GQA [16], covering object grounding and relational reasoning. These examples illustrate the model’s of spatial understanding and positional reasoning, which is crucial for language-guided UAV navigation.
Figure 10: Qualitative results on trajectory summary task. The model summarizes the route by identifying key landmarks (building, pipe, walkway, garage) and the major motion segments (take-off, left turns, forward movement), demonstrating temporal understanding of the visual trajectory.
IV-B4 Qualitative Analysis
We provide qualitative navigation trajectories in Fig. 8 to assess the model’s ability to interpret and ground natural language instructions in realistic outdoor environments. The first example illustrates a medium-length instruction including several landmark-based references. Our agent demonstrates reliable instruction-following behavior by consistently grounding entities such as shops, bike parking, and green shop, and by executing the corresponding spatial maneuvers (e.g., turning, approaching, passing) at the appropriate visual cues. This indicates that the model preserves strong cross-modal alignment between linguistic expressions and visual observations, a fundamental capability for step-by-step decision-making in VLN. The second example showcases navigation under a significantly longer and more descriptive instruction spanning over one hundred actions. Throughout this extended trajectory, the agent maintains instruction consistency, preserves stable temporal grounding, and generates coherent movement decisions despite the increased complexity. These results highlight the robustness of our model in long-horizon instruction following, a key requirement for practical UAV VLN in unconstrained outdoor environments. To analyze the trajectory behavior of our navigation policy,
Fig. 9 presents qualitative examples from the auxiliary VQA datasets used in training. On Open3D-VQA dataset, the model correctly answers 3D spatial queries such as which object is higher or what lies to the left/right in an egocentric aerial view. On the GQA subset, it identifies objects and resolves simple spatial relations in visually cluttered scenes. These examples show that the model can interpret scene geometry and spatial layout from egocentric inputs. Fig. 10 presents qualitative examples from the trajectory reasoning task. Given the historical egocentric visual trajectory, the model produces a concise natural-language description that captures the high-level motion pattern and the principal scene transitions along the route. The generated summaries demonstrate consistent temporal abstraction by identifying turning segments, periods of forward movement, and transitions associated with spatial landmarks such as parks, intersections, and surrounding buildings. Moreover, the model reflects the causal structure of the trajectory by linking sequences of actions to the corresponding visual and semantic changes in the environment. These examples show that the model can integrate long-term dependencies and organize extended action sequences into coherent semantic units, which provides a complementary form of temporal grounding beyond low-level action prediction.
TABLE V: Comparison of Model Efficiency.
Method Peak GPU Memory Inference Latency
OpenFly (7B) $\sim$ 14 GB 0.5 s
Ours (8B) $\sim$ 17 GB 0.7 s
Ours (8B)-AWQ $\sim$ 6 GB 0.5 s
IV-C Model Efficiency
We evaluate the inference efficiency of our method to assess its practicality for aerial vision-language navigation. We compare our approach with OpenFly [12], a recent state-of-the-art aerial VLN method, as both methods adopt a VLM-based design paradigm. All measurements are conducted on a single NVIDIA A100 GPU. We report end-to-end inference latency per step  [^1] End-to-end inference latency is measured as the wall-clock time from receiving multimodal observations (RGB images and language instruction) to producing the predicted action. together with peak GPU memory. As shown in Table V, the full-precision 8B model exhibits a per-step inference latency of about 0.7 s and a peak GPU memory usage of approximately 17 GB. We further note that the favorable inference latency of OpenFly [12] mainly benefits from its limited visual input, leading to lower multimodal processing overhead and reduced temporal visual context. To alleviate the memory overhead, we further evaluate an Activation-aware Weight Quantization (AWQ) [22] version of the 8B model. AWQ substantially reduces the memory footprint and lowers the inference latency to around 0.5 s, resulting in a decision rate close to 2 Hz. Further improvements in inference efficiency are expected to benefit from more advanced acceleration techniques and hardware-aware optimizations.
IV-D Ablation Study
TABLE VI: Ablation study on auxiliary tasks: Spatial Perception (SP) and Trajectory Reasoning (TR). Both tasks improve success rate (SR) and trajectory quality (SDTW) in both seen and unseen environments.
Auxiliary Tasks Val Seen Val Unseen
SP TR $uparrow SR\uparrow SDTW\uparrow$
9.6 4.5 5.8 1.6
10.8 5.4 7.0 1.7
11.4 6.3 8.1 2.2
IV-D1 Effect of Auxiliary Tasks
We conduct ablation experiments to evaluate the effect of the two auxiliary tasks: Spatial Perception (SP) and Trajectory Reasoning (TR) in Table VI. Without auxiliary supervision, the policy shows limited spatial grounding and produces the lowest SR and SDTW on both splits. SP leads to consistent gains, confirming that geometric cues improve local action grounding. Incorporating TR further boosts performance, with the largest improvements on unseen environments, where long-horizon cues are critical for handling novel layouts. These results demonstrate that the two objectives are complementary: SP enhances local spatial understanding, while TR strengthens temporal reasoning. Their combination yields a more reliable navigation policy and supports the design of our multi-task framework, reinforcing our goal of unifying spatial, temporal, and embodied reasoning within a single model.
IV-D2 Effect of History Representation Strategies
TABLE VII: Comparison of different history representation strategies on AerialVLN-S Val-Unseen split.
[…]
Current-Only $(2B)$ 4.1 1.1
Short-Term Memory Bank $(2B,k=8)$ 4.3 1.2
Long-Horizon Uniform Sampling $(2B,k=8)$ 5.5 1.6
Long-Horizon Uniform Sampling $(8B,k=8)$ 5.9 1.6
Figure 11: Success rate across different trajectory lengths on the AerialVLN-S Val-Unseen split. Long-Horizon Uniform Sampling maintains stable performance across episodes of varying lengths.
We compare three history representation strategies. (1) Current-Only, where the model relies solely on the latest observation, often leading to unstable behaviors such as repeated fine-grained heading corrections. (2) Short-Term Memory Bank, which maintains a fixed-size memory bank following a first-in-first-out (FIFO) policy, keeping only the most recent observations and discarding earlier ones [12]. This provides short-range temporal cues while discards earlier information essential for global navigation. (3) Long-Horizon Uniform Sampling, which uniformly samples a small set of frames from the entire past trajectory to provide long-range visual context. Table VII compares these history representation strategies. Among them, Long-Horizon Uniform Sampling ($k=8$) achieves the best overall performance. By drawing observations uniformly from entire past trajectory, this strategy preserves both short-term motion cues and long-range scene information. As a result, it provides a more reliable temporal context than relying solely on the current frame or restricting history to a short FIFO queue. We compares these strategies using the NVILA-2B model for computational efficiency. The larger 8B variant follows the same trend and achieves consistently higher accuracy.
To further analyze how different history strategies behave under varying temporal horizons, we group the Val-Unseen episodes into three categories based on their trajectory length and report the per-step success rate for each method in Fig. 11. The Current-Only baseline performs reasonably well on short trajectories because the goal is often visible early, but its accuracy declines sharply as the trajectory becomes longer, indicating limited robustness without historical context. The Long-Horizon Uniform Sampling strategy maintains higher performance across all trajectory lengths, with clear advantages on medium and long episodes where global scene information becomes crucial. These results highlight the importance of accessing long-range temporal cues for stable navigation in large, visually diverse environments.
TABLE VIII: Ablation study on the effect of action merging (AM) , keyframe selection (KS) and label reweighting (LR) on AerialVLN-S Val-Seen split. All of them consistently improve performance across all metrics.
[…]
90.2 1.8 20.5 0.7
91.5 3.0 26.8 1.4
87.1 5.7 28.9 1.9
82.1 8.1 32.2 3.8
IV-D3 Effect of Action Merging, Keyframe Selection, and Label Reweighting
Table VIII reports the effect of action merging (AM), keyframe selection (KS), and label reweighting (LR). For computational efficiency, this ablation uses the NVILA-2B model without auxiliary tasks, and all variants adopt a long-horizon uniform sampling of 8 history frames. Enabling action merging leads to immediate gains, as consolidating micro-actions into coherent motion segments produces cleaner supervisory signals compared to raw steps. Adding keyframe selection further improves navigation accuracy. Selecting the boundaries of merged actions as keyframes provides a more structured trajectory history in which visual transitions align with meaningful motion changes. This reduces temporal redundancy while preserving observations most relevant to spatial landmarks. Finally, incorporating label reweighting achieves the best overall performance. By balancing the contribution of long merged segments and short corrective actions, it prevents the model from overfitting to frequent yet trivial motions, thereby strengthening action-level supervision.
We further ablate the maximum number of consecutive actions merged (i.e., the merge step) and report the results in Table IX. As shown, a merge step of 3 achieves the best performance, while a merge step of 1 (i.e., no merging) retains the original fragmented action sequence and leads to lower navigation performance. Excessively large merge steps (e.g., 6) slightly degrade performance, likely because they skip important visual cues and reduce the model’s ability to adapt to local changes. We therefore set the maximum merge length to 3. Keyframes are then selected at the boundaries of these refined segments, ensuring sufficient temporal coverage and preserving landmarks that may appear along extended straight paths. Overall, the combined use of AM, KS, and LR yields substantial improvements in SR and SDTW, demonstrating that structuring the action sequence and rebalancing its supervision are both crucial for reliable aerial navigation.
TABLE IX: Ablation on the maximum number of consecutive actions merged evaluated on AerialVLN‑S val‑seen split. The best performance is achieved with a merge step of 3.
[…]
1 90.2 1.8 20.5 0.7
3 87.1 5.7 28.9 1.9
6 89.5 5.1 25.3 1.4
TABLE X: Quantitative failure case analysis on the AerialVLN-S val-unseen split. Note that $\dagger$ indicates our method without auxiliary tasks. $\Delta$ denotes the difference in episode counts between the proposed model and the variant trained without auxiliary tasks.
Failure Type Ours Ours $\Delta$
Perception-related 194 204 10-10
Long-horizon drift 124 135 11-11
Collision 142 150 8-8
Stop failure 107 90 +17+17
IV-E Failure Case Analysis
Although our method achieves strong performance and outperforms competitive baselines, it is still susceptible to several typical failure modes. To better understand the remaining challenges and guide future improvements, we further analyze its most common failure cases below. A common failure mode arises from ambiguous instructions that provide only a sequence of low-level actions without any accompanying landmark references (e.g., “rotate, go straight, turn left, turn right, hold on”). The absence of such grounding references makes it difficult for the agent to determine where in the environment each action should be executed, leading to misalignment between the instruction and the visual context. We manually identify 21 severely underspecified instructions whose action sequences lack explicit landmark references. On this subset, both the proposed model and the variant without auxiliary tasks fail. This observation is consistent with previous work[45] as the absence of explicit landmark references prevents the agent from grounding the action sequence to specific locations in the environment, leading to inherently unreliable decision-making. Excluding these cases, we analyze the remaining episodes using trajectory-level metrics and conduct a quantitative analysis on the AerialVLN-S val-unseen split. Failure cases are categorized into stop failures (OSR=1 with incorrect termination), collision failures (OSR=0 with obstacle contact), long-horizon drift (OSR=0 with low trajectory similarity measured by nDTW), and perception-related failures (remaining cases exhibiting partial trajectory alignment but incorrect intermediate grounding). As summarized in Table X, the auxiliary tasks reduce both long-horizon drift and perception-related failures compared to the model trained without auxiliary supervision. This suggests improved global alignment over extended trajectories and more stable visual grounding under monocular observation. Collision failures are slightly reduced. In contrast, stop failures increase, indicating that the auxiliary tasks enable the agent to reach the goal region more frequently while termination decisions remain imperfect. Overall, the results show that the auxiliary tasks mitigate specific failure modes, particularly long-horizon drift and perception-related errors, by strengthening alignment between visual perception, textual language, and embodied action.
V Conclusion and Discussion
In this paper, we present a unified framework for aerial vision-and-language navigation that operates solely on monocular egocentric RGB inputs and natural language instructions, eliminating the need for panoramic observations, depth sensing, or external localization modules. To address core challenges posed by the complexity of large-scale outdoor scenes and the long-horizon nature of aerial navigation, we reformulate aerial VLN as a next-token prediction problem and integrates spatial perception, trajectory reasoning, and action generation through task-specific prompts. Experiments on the Aerial VLN benchmark demonstrate that our method achieves new state-of-the-art performance among RGB-only approaches and significantly narrows the gap with panoramic RGB-D counterparts. This work highlights the potential of scalable, prompt-driven multimodal learning as an efficient solution for deploying embodied aerial agents in real-world scenarios.
Despite the encouraging results, several limitations remain. First, the current framework relies on passive monocular observations, which may lead to incomplete environmental coverage under restricted field-of-view settings. Second, performance can degrade under ambiguous instructions that lack explicit landmark references. Third, long-horizon trajectory alignment remains challenging when small prediction errors accumulate over time. Future work will explore active viewpoint control to expand perceptual coverage, ambiguity-aware instruction clarification to improve grounding under underspecified commands, and hierarchical or staged planning strategies to enhance long-range consistency. Moreover, aerial VLN method could benefit from lightweight obstacle-aware planning modules that operate alongside the navigation policy.